# Heterogeneous Position Effects and the Power of Rankings

### 21 Dec 2023 v3

@@ Abstract
Many online retailers and search intermediaries present products on ranked product lists. Because changing the ordering of products influences their revenues and consumer welfare, these platforms can design rankings to maximize either of these two metrics. In this paper, I study how rankings following these two objectives differ and what determines this difference. First, I highlight that when position effects are heterogeneous, rankings can increase both revenues and consumer welfare by increasing the overall conversion rate. Second, I provide empirical evidence for this heterogeneity, highlighting that cheaper alternatives have stronger position effects. Finally, I quantify revenue and consumer welfare effects across the different rankings. To this end, I develop an estimation procedure for the search and discovery model of Greminger (2022) that yields a smooth likelihood function by construction. The results from the model show that rankings designed to maximize revenues can also increase consumer welfare. Moreover, these revenue-based rankings reduce consumer welfare only to a limited extent relative to a consumer-welfare-maximizing ranking.
@@ Introduction
In this paper, I ask whether this is the case and study how rankings following the two objectives differ in how they affect consumers and a platform’s revenues from a price-based commission. First, I use a simple example to highlight a crucial factor determining such differences: heterogeneous position effects. If there are position effects, a ranking can shift demand to more expensive alternatives by moving them higher on the list. However, suppose position effects are heterogeneous and cheaper alternatives have stronger position effects. In this case, moving expensive alternatives up on the list also decreases the conversion rate such that maximizing the platform’s revenues requires balancing the two effects. Hence, depending on whether position effects are heterogeneous, a revenue-maximizing ranking will move more or less expensive alternatives to the top and affect consumers accordingly.
Whether cheaper alternatives have stronger position effects such that a platform needs to balance the two effects is an empirical question and not clear a priori. To address it, I use experimental ranking variation to identify heterogeneous position effects in data from a large online search intermediary, Expedia. The results show that cheaper alternatives indeed have stronger position effects. More broadly, I find that product attributes and the position on the list complement each other such that hotels with desirable attributes have stronger position effects. Hence, moving hotels with a lower price and other desirable attributes higher up on the list will increase the conversion rate on average, potentially increasing both revenues and consumer welfare.
Analyzing whether this is the case requires quantifying how the different rankings affect platform revenues and consumer welfare. To this end, I build on and estimate the search and discovery (SD) model of Greminger (2022). In the model, consumers decide between discovering more products by scrolling down the list and learning more about an already-discovered product by going to its detail page. I argue that this model is well-suited to study the consumer welfare effects of different rankings because it provides a micro-foundation for position effects: consumers decide to stop scrolling before ever reaching the lower positions on the list. Moreover, I show that this discovery mechanism implies position effects that closely match the pattern revealed by the descriptive evidence: alternatives with desirable attributes have stronger position effects. As a result, the estimated model predicts heterogeneous position effects that closely match those in the data and, hence, fits the moments determining the trade-off between the different rankings notably well.
@@
To address the substantive research question, I also make several methodological contributions. First, I develop an estimation approach for the SD model that yields a smooth likelihood by construction and avoids computationally expensive steps during estimation. Second, I extend the analysis of Greminger (2022) to show how the SD model captures position effects conditional on the observed product attributes. Finally, I use this result to show how the observed position effects are sufficient to identify model parameters without observing how many products consumers discover.
@@
This paper contributes to the literature examining the effects of different rankings on consumer welfare, search behavior, and platform revenues. Initial studies in this literature focus on the effects of utility-based rankings (Ghose et al., 2012, 2014; Ursu, 2018). More recent studies additionally consider the effects of rankings designed to increase platform revenues, documenting possible trade-offs between making consumers better off and increasing platform revenues (Choi and Mela, 2019; Zhang et al., 2021; Compiani et al., 2023; Donnelly et al., 2023). I add to this literature by highlighting how this trade-off stems from heterogeneous position effects, deriving how these heterogeneous position effects arise in a micro-founded model, and quantifying this trade-off by estimating the same model. By focusing on the effects of rankings with different objectives, this paper also differs from recent work by Zhang et al. (2023), who estimate the SD model to study how different search routes affect consumers’ search behavior.
@@
By providing empirical evidence that shows how position effects depend on product attributes like price, I also add to the empirical literature that documents position effects on ranked product lists (Ghose et al., 2012, 2014; De Los Santos and Koulayev, 2017; Ursu, 2018) and in search advertising (Goldman and Rao, 2014; Athey and Imbens, 2015; Jeziorski and Segal, 2015; Jeziorski and Moorthy, 2018). [^1] These studies find that ads with a larger click-through-rate on any position have weaker position effects. The difference to my results suggests that consumers use a general-purpose search engine like Google with a specific outcome already in mind instead of using the list to compare different alternatives. This is corroborated by results of Blake et al. (2015), who find that search advertising is not effective for a well-known platform like eBay. Unlike Ghose et al. (2014), who also consider how position effects vary with product attributes, I use experimental variation in rankings to account for endogeneity in rankings and consistently find that desirable attributes amplify position effects.
@@
The methodological contributions in this paper add to the empirical search literature that has used a range of alternative models to rationalize position effects (e.g., Chan and Park, 2015; Chen and Yao, 2017; De Los Santos and Koulayev, 2017; Ursu, 2018; Choi and Mela, 2019; Gibbard, 2022b; Compiani et al., 2023; Donnelly et al., 2023). The SD model I estimate has several features that make it particularly useful to capture position effects and quantify the effects of different rankings, and I highlight these features and how they relate to other sequential search models after introducing the model. As they are nested in the SD model, my estimation approach can also be used to estimate directed search models á la Weitzman (1979) and, hence, adds to the recent literature introducing estimation approaches without researcher-chosen smoothing parameters for these models Jiang et al. (2021); Chung et al. (2023). However, unlike these approaches, my approach does not require observing the search order, is easier to implement, and requires fewer computations. [^2] The approaches of Jiang et al. (2021) and Chung et al. (2023) grow in computational complexity in the number of clicks each consumer makes, which is not the case with my approach.
@@
By showing that a utility-based ranking maximizes consumer welfare in the SD model and introducing two heuristics for maximizing revenues in the short-term, I also add to the literature deriving algorithms to optimally choose in which order to display alternatives to consumers (e.g., Ryzin and Mahajan, 1999; Talluri and Van Ryzin, 2004; Davis et al., 2014; Ursu and Dzyabura, 2020; Chu et al., 2020; Derakhshan et al., 2022; Compiani et al., 2023). Whereas this literature motivates this paper’s empirical evaluation of how revenue-maximizing rankings affect consumers, it tends to rely on more stylized demand models and, [^3] Early work in this literature did not consider search. The model of Chu et al. (2020) precludes position effects for consumers who do not buy an alternative, contrasting what I observe in the data. Derakhshan et al. (2022) consider heterogeneous position effects in the context of maximizing the conversion rate (as opposed to revenues) and assume that consumers first decide how many products to discover before inspecting all of the discovered products at the same time. This model also contrasts the data as almost all consumers skip clicking on some alternatives. Ursu and Dzyabura (2020) instead focus on a retailer deciding how to place independent categories within a retailer. except for Compiani et al. (2023), does not quantify the different effects.
@@
I use click-stream data from a leading online travel agency, Expedia. The data can be obtained from Kaggle.com and contains information on clicks and purchases for 166,039 search sessions between November 2012 and June 2013.555The data is available under the following link: https://www.kaggle.com/c/expedia-personalized-sort/data. A search session starts at the point where a consumer has submitted a query for a hotel stay on Expedia. Following this query, Expedia presents a list of the available options on which consumers observe various hotel characteristics, such as the price per night and the review score. They then can scroll down to reveal more hotels on the list or click on a particular item they already see on the list. Clicking on an item leads to the hotel’s detail page, which reveals further information and allows booking the hotel. Ursu (2018) provides a more comprehensive discussion of the dataset, and I apply similar criteria to prepare the final sample (see Appendix E).

### 29 Dec 2023 v4

@@ Abstract
Many online retailers and search intermediaries present products on ranked product lists. Because changing the ordering of products influences their revenues and consumer welfare, these platforms can design rankings to maximize either of these two metrics. In this paper, I study how rankings serving these two objectives differ and what determines this difference. First, I highlight that when position effects are heterogeneous, rankings can increase both revenues and consumer welfare by increasing the overall conversion rate. Second, I provide empirical evidence for this heterogeneity, highlighting that cheaper alternatives have stronger position effects. Finally, I quantify revenue and consumer welfare effects across the different rankings. To this end, I develop an estimation procedure for the search and discovery model of Greminger (2022) that yields a smooth likelihood function by construction. The results from the model show that rankings designed to maximize revenues can also increase consumer welfare. Moreover, these revenue-based rankings reduce consumer welfare only to a limited extent relative to a consumer-welfare-maximizing ranking.
@@ Introduction
In this paper, I ask whether this is the case and study how rankings serving the two objectives differ in their impact on consumers and the platform’s revenues. First, I use a simple example to highlight a crucial factor determining such differences: heterogeneous position effects. If a platform takes a price-based commission or markup, shifting demand to more expensive alternatives will increase the platform’s revenues. Hence, a ranking can increase the platform’s revenues by exploiting position effects and moving more expensive alternatives higher up on the list. However, suppose position effects are heterogeneous and cheaper alternatives have stronger position effects. In this case, moving expensive alternatives up on the list also decreases the conversion rate. As a result, maximizing the platform’s revenues requires balancing the two effects.
Whether position effects are heterogeneous such that a platform needs to balance the two effects is an empirical question and not clear a priori. To address it, I use experimental ranking variation to identify heterogeneous position effects in data from a large online search intermediary. The results show that cheaper alternatives indeed have stronger position effects. More broadly, I find that product attributes and the position on the list complement each other such that alternatives with desirable attributes have stronger position effects. Hence, moving products with a lower price and other desirable attributes higher up on the list will increase the conversion rate on average, potentially increasing both revenues and consumer welfare.
Analyzing whether this will occur requires quantifying how the different rankings affect the platform’s revenues and consumer welfare. To this end, I build on and estimate the search and discovery (SD) model of Greminger (2022). In the model, consumers decide between discovering more products by scrolling down the list and learning more about an already-discovered product by going to its detail page. I argue that this model is well-suited to study the consumer welfare effects of different rankings because it provides a micro-foundation for position effects: consumers decide to stop scrolling before ever reaching the lower positions on the list. Moreover, I show that this discovery mechanism implies position effects that closely match the pattern revealed by the descriptive evidence. As a result, the estimated model predicts heterogeneous position effects that closely match those in the data and, hence, fits the moments determining the trade-off between the different rankings notably well.
@@
To address the substantive research question, I also make several methodological contributions. First, I develop an estimation approach for the SD model that yields a smooth likelihood by construction and avoids computationally expensive steps during estimation. Second, I extend the analysis of Greminger (2022) to show how the SD model captures position effects conditional on the observed product attributes. Finally, I use this result to show how the observed position effects are sufficient to identify the model parameters without observing how many products consumers discover.
@@
This paper contributes to the literature examining the effects of different rankings on consumer welfare, search behavior, and platform revenues. Several studies in this literature focus on the effects of utility-based rankings (Ghose et al., 2012, 2014; Ursu, 2018). More recent studies additionally consider the effects of rankings designed to increase platform revenues, documenting possible trade-offs between making consumers better off and increasing platform revenues (Choi and Mela, 2019; Zhang et al., 2021; Compiani et al., 2023; Donnelly et al., 2023). I add to this literature by highlighting how this trade-off stems from heterogeneous position effects, deriving how these heterogeneous position effects arise in a micro-founded model, [^1] While heterogeneous position effects may also arise in other empirical search models, prior work has not shown whether this is the case and how these alternative models generate heterogeneous position effects. Hence, it remains unclear whether these models would also fit these crucial moments in the data. and quantifying this trade-off by estimating the same model. By focusing on the effects of rankings, this paper also differs from recent work by Zhang et al. (2023), who estimate the SD model to study how different search routes affect consumers’ search behavior.
@@
By providing empirical evidence that shows how position effects depend on product attributes like price, I also add to the empirical literature that documents position effects on ranked product lists (Ghose et al., 2012, 2014; De Los Santos and Koulayev, 2017; Ursu, 2018) and in search advertising (Goldman and Rao, 2014; Athey and Imbens, 2015; Jeziorski and Segal, 2015; Jeziorski and Moorthy, 2018). [^2] These studies find that ads with a larger click-through rate on any position have weaker position effects. The difference to my results suggests that consumers use a general-purpose search engine like Google with a specific outcome already in mind instead of using the list to compare different alternatives. This is corroborated by results of Blake et al. (2015), who find that search advertising is not effective for a well-known platform like eBay. Unlike Ghose et al. (2014), who also consider how position effects vary with product attributes, I use experimental variation in rankings to account for endogeneity in rankings and consistently find that desirable attributes amplify position effects.
@@
The methodological contributions in this paper add to the empirical search literature that has used a range of alternative models to rationalize position effects (e.g., Chan and Park, 2015; Chen and Yao, 2017; De Los Santos and Koulayev, 2017; Ursu, 2018; Choi and Mela, 2019; Gibbard, 2022b; Compiani et al., 2023; Donnelly et al., 2023). The SD model has several features that make it particularly useful to capture position effects and quantify the effects of different rankings. I highlight these features and relate them to other sequential search models after introducing the model. As they are nested in the SD model, my estimation approach can also be used to estimate directed search models á la Weitzman (1979) and, hence, adds to the recent literature introducing estimation approaches without researcher-chosen smoothing parameters for these models Jiang et al. (2021); Chung et al. (2023). However, unlike these approaches, my approach does not require observing the search order and requires fewer computations. [^3] The approaches of Jiang et al. (2021) and Chung et al. (2023) grow in computational complexity in the number of clicks each consumer makes, which is not the case with my approach.
@@
By showing that a utility-based ranking maximizes consumer welfare in the SD model and introducing two heuristics for maximizing revenues in the short-term, I also add to the literature deriving algorithms to optimally choose in which order to display alternatives to consumers (e.g., Ryzin and Mahajan, 1999; Talluri and Van Ryzin, 2004; Davis et al., 2014; Ursu and Dzyabura, 2020; Chu et al., 2020; Derakhshan et al., 2022; Compiani et al., 2023). Whereas this literature motivates this paper’s empirical evaluation of how revenue-maximizing rankings affect consumers, it tends to rely on more stylized demand models and, [^4] Early work in this literature did not consider search. The model of Chu et al. (2020) precludes position effects for consumers who do not buy an alternative, contrasting what I observe in the data. Derakhshan et al. (2022) consider heterogeneous position effects in the context of maximizing the conversion rate (as opposed to revenues). Their model implies that consumers inspect all of the alternatives on the product list up to some rank, which contrasts the data as almost all consumers skip clicking on some alternatives. Ursu and Dzyabura (2020) instead focus on a retailer deciding how to place independent categories within a retailer. except for Compiani et al. (2023), does not quantify the different effects.
@@
I use click-stream data from a leading online travel agency, Expedia. The data can be obtained from Kaggle.com and contains information on clicks and purchases for 166,039 search sessions between November 2012 and June 2013.666The data is available under the following link: https://www.kaggle.com/c/expedia-personalized-sort/data. A search session starts at the point where a consumer has submitted a query for a hotel stay on Expedia. Following this query, Expedia presents a list of the available options on which consumers observe various hotel characteristics, such as the price per night and the review score. They then can scroll down to reveal more hotels on the list or click on a particular item they already see on the list. Clicking on an item leads to the hotel’s detail page, which reveals further information and allows booking the hotel. Ursu (2018) provides a more comprehensive discussion of the dataset, and I apply similar criteria to prepare the final sample (see Appendix E).